OpenStudy (hellomynameiskam):

What is the difference? - 1/10 - 8/10 A. - 9/10 B. - 7/10 C. 7/10 D. 9/10 @rebeccaxhawaii

OpenStudy (anonymous):

maybe I can help?

rebeccaxhawaii (rebeccaxhawaii):

add the two

rebeccaxhawaii (rebeccaxhawaii):

add them and they will remain negative

rebeccaxhawaii (rebeccaxhawaii):

what do you get ?

OpenStudy (hellomynameiskam):

- 9/10

OpenStudy (anonymous):

no

rebeccaxhawaii (rebeccaxhawaii):

yes that is correct

OpenStudy (alexandervonhumboldt2):

\[-\frac{ 1 }{ 10 }-\frac{ 8 }{ 10 }=-\frac{ (1+8) }{ 10 }\]

OpenStudy (alexandervonhumboldt2):

yes -9/10 is correct

OpenStudy (anonymous):

oops never mind